# Stackelberg-Pareto Synthesis with Quantitative Reachability Objectives

### v2

@@ 4.3 Bounding the Minimum Component of Pareto-Optimal Costs
[…]
The proof of Lemma 4.6 requires the next technical property about trees. We recall the notion of depth of a node in a tree: the root has depth 00, and if a node has depth $d$, then its sons have depth $d+1$.
@@
When $n=1$, the worst case is the tree composed of one branch with $\ell-1$ nodes. When $n\geq 2$, the worst case happens when the tree is binary, all its leaves have the same depth, and along any branch, there are $\ell-1$ consecutive nodes with degree 1 before and after each node with degree 2. In both cases $n=1$ and $n\geq 2$, the worse tree has $n=2^{k}$ leaves for some $k\geq 0$. We denote this tree by $\mathcal{T}_{k}$ (see Figure 4).
Figure 4: Tree $\mathcal{T}_{k}$ with $k=2$ and $\ell=3$
Let us prove by induction on $k$ that the depth of the leaves of $\mathcal{T}_{k}$ is bounded by $\ell\cdot(\log_{2}(n)+1)=\ell\cdot(k+1)$. When $k=0$, the unique leaf of $\mathcal{T}_{0}$ has depth $\ell-2\leq\ell$. When $k\geq 1$, consider the node $v$ of $\mathcal{T}_{k}$ with degree 2 and with the smallest depth $d$. The nodes with depth $<d$ are ancestors of $v$ and there are $\ell-1$ of them. Hence $d=\ell-1$. Moreover, the two subtrees of $v$ are isomorphic to $\mathcal{T}_{k-1}$. Therefore, by induction hypothesis, the leaves of $\mathcal{T}_{k}$ have depth bounded by $(\ell-1)+1+\ell\cdot k\leq\ell\cdot(k+1)$.
@@
Let us construct the first Pareto-optimal cost $c^{(0)}$. Let $h_{0}$ be the history of maximal length such that $h_{0}\rho$ and $w(h_{0})=B$. This history $h_{0}$ exists because the arena is binary and $\textsf{cost}({\rho})=d$ with $d_{min}$ satisfying (12). Moreover, as $\sigma_{0}$ is a solution, $h_{0}$ visits Player 00’s target. We consider the subtree $\mathcal{T}_{0}$ of $\mathcal{T}$ rooted in the last vertex of $h_{0}$. This subtree has at most $|C_{\sigma_{0}}|$ leaves as it is the case for $\mathcal{T}$. Its nodes with degree $\geq 2$ correspond to histories that are branching points [^9] We recall that a branching point is a history which is the greatest common prefix of two witnesses.. Moreover, any two consecutive nodes with degree one along any branch of ${\mathcal{T}}_{0}$ are in the same region, as Player 00’s target is visited by $h_{0}$ and the first visit to a target of Player $1$ is in a leaf of $\mathcal{T}_{0}$. As $\sigma_{0}$ is without cycle in the sense of Lemma 3.3, it follows that there are at most $|V|-1$ consecutive nodes with degree one along any branch of $\mathcal{T}_{0}$.
@@
Let us construct the second Pareto-optimal cost $c^{(1)}$ as we did for $c^{(0)}$. Let $h_{1}$ be the history of maximal length such that $h_{1}\rho$ and $w(h_{1})=B+\delta+1+f({0},{t})$ (notice that we use the bound of (14)). This history $h_{1}$ exists for the same reasons as for $h_{0}$. Let $\mathcal{T}_{1}$ be the subtree rooted in the last vertex of $h_{1}$ (it is a subtree of $\mathcal{T}_{0}$). We can apply Lemma 4.7 as for $\mathcal{T}_{0}$: the leaves of $\mathcal{T}_{1}$ with the smallest depth have a depth $\leq\delta$ Thus, we set $c^{(1)}$ as the cost of a witness associated with one of these leaves. We get that $c^{(1)}_{min}\leq B+\delta+1+f({0},{t})+\delta$ by definition of $w(h_{1})$. By (7), we get
@@
Finally, we gather all our intermediate results and combine them, in a way to complete the proof of Theorem 4.1. Thanks to Lemmas 4.2-4.6, calculations can be done in a way to have an explicit formula for $f({B},{t})$ and a bound on its value.
@@
It follows that $f({B},{t+1})$ can be computed thanks to $f({0},{t})$ bounded by induction:
@@
Therefore, Inequation (4) holds for $f({B},{t+1})$. This completes the proof of Theorem 4.1 by induction on $t$.
@@
In the sequel, we use the same notation $f({B},{t})$ for any SP games, having in mind that $|V|$ has to be multiplied by $W$ when the game arena is not binary.
@@ 5.1 Finite-Memory Solutions
  • We show that at any deviation [^11] We recall that a deviation is a history $hv$ with $h\in\textsf{Hist}^{1}$, $v\in V$, such that $h$ is prefix of some witness, but $hv$ is prefix of no witness., Player 00 can switch to a punishing strategy that imposes that the consistent plays $\pi$ either satisfy $\textsf{val}(\pi)\leq B$ or $\textsf{cost}({\pi})$ is not Pareto-optimal. Moreover, this punishing strategy is finite-memory with an exponential memory.
@@ 5.2 Punishing Strategies
Let $\sigma_{0}$ be a Pareto-bounded solution to the SPS Problem. By Theorem 4.1 (3) (4), we get that $c_{i}\leq f({B},{t})$ or $c_{i}=\infty$ for all $c\in C_{\sigma_{0}}$ and all $i\in\{1,\ldots,t\}$, such that $f({B},{t})$ is exponentially bounded. Moreover, if a play $\rho$ is Pareto-optimal, then $\textsf{val}(\rho)\leq B$. We define for each history $g\in\textsf{Hist}_{\sigma_{0}}$ its record $\textsf{rec}({h})=$ $(w(h),\textsf{val}(h),\textsf{cost}({h}))$ whose values are truncated to $\infty$ if they are greater than $f({B},{t})$.
@@
Let $hv$ be a deviation such that $\textsf{val}(h)=\infty$, that is, $h$ does not visit Player 00’s target. As $\sigma_{0}$ is a solution, notice that $\sigma_{0|hv}$ imposes to each consistent play $\pi$ in $G_{|hv}$ to satisfy $\textsf{val}(\pi)\leq B$ or $\textsf{cost}({\pi})>c$ for some $c\in C_{\sigma_{0}}$. We are going to replace $\sigma_{0|hv}$ by a winning strategy in a zero-sum game $H$ with an exponential arena and an omega-regular objective [^12] We assume that the reader is familiar with the concept of zero-sum game with an omega-regular objective, see e.g. [25]. that is equivalent to what $\sigma_{0|hv}$ imposes to plays. The arena of $H$ is the arena of $G$ extended with the record $\textsf{rec}({g})$ of the current history $g$. More precisely, vertices are of the form $(v,(m_{1},m_{2},m_{3}))$ with $v\in V$, $m_{1},m_{2}\in\{0,\ldots,f({B},{t})\}\cup\{\infty\}$, and $m_{3}\in\big{(}\{0,\ldots,f({B},{t})\}\cup\{\infty\}\big{)}^{t}$, such that, whenever $v$ belongs to some target, the weight component $m_{1}$ allows to update the (truncated) val component $m_{2}$ and the (truncated) cost component $m_{3}$. The initial vertex of $H$ is equal to $(v,\textsf{rec}({hv}))$. This arena is finite and of exponential size by the way the values of $(m_{1},m_{2},m_{3})$ are truncated and by Theorem 4.1 (4). The omega regular objective of $H$ is the disjunction between
@@
Suppose now that $hv$ is a deviation with $\textsf{val}(h)<\infty$. As $h$ already visits Player 00’s target, we can use, in place of $\sigma_{0|hv}$, any memoryless strategy $\tau^{\textsf{Pun}}_{v,\textsf{rec}({hv})}$ as deviating strategy.
@@ 5.3 Lasso Witnesses
(1) We first suppose that $C_{\sigma_{0}}\neq\{(\infty,\ldots,\infty)\}$. According to Theorem 4.1, from $\sigma_{0}\in\mbox{SPS}(G,B)$, one can construct a solution $\sigma_{0}^{\prime}\in\mbox{SPS}(G,B)$ that is Pareto-bounded, thus without cycles. Consider a set of witnesses $\textsf{Wit}_{\sigma^{\prime}_{0}}$ for this strategy, and the region decomposition of its witnesses. Along any witness $\rho$, according to Lemma 3.3, it is thus impossible to eliminate cycles. Recall that Lemma 3.3 only considers cycles implying histories in the same region, those cycles with a null weight before visiting Player 00’s target; moreover, the last region of $\rho$ is excluded. Let us study the memory used by $\sigma^{\prime}_{0}$. Along a witness $\rho$:
  • As $\textsf{val}(\rho)\leq B$ and there is no cycle with a null weight before visiting Player 00’s target, it follows that the smallest prefix $h$ of $\rho$ such that $\textsf{val}(h)=\textsf{val}(\rho)$ has a length $|h|\leq|V|\cdot B\cdot W$ (the presence of $W$ comes from the fact that the arena has been made binary).
  • Once Player 00’s target is visited and before the last region, $\sigma^{\prime}_{0}$ is “locally” memoryless inside each region, as there is no cycle.
@@
  • The number of regions traversed by a witness $\rho$ is bounded by $(t+2)\cdot|\textsf{Wit}_{\sigma^{\prime}_{0}}|$. Indeed, the number of visited targets increases from 00 to $t+1$ and the set $\textsf{Wit}_{\sigma^{\prime}_{0}}(h)$ decreases until being equal to $\{\rho\}$. As there are $|C_{\sigma^{\prime}_{0}}|$ witnesses, the total number of traversed regions is bounded by $(t+2)\cdot|C_{\sigma^{\prime}_{0}}|^{2}$ which is exponential by Corollary 4.11.
@@
(2) Suppose that $C_{\sigma_{0}}=\{(\infty,\ldots,\infty)\}$. This means that each play $\rho\in\textsf{Play}_{\sigma_{0}}$ visits no target of Player $1$ and that $\sigma_{0}$ can impose to each such $\rho$ to have a value $\textsf{val}(\rho)\leq B$. As done in the proof of Lemma 5.2, we can replace $\sigma_{0}$ by a strategy corresponding to a winning strategy in a zero-sum game $H$ with an exponential arena and a reachability objective. The arena of $H$ has vertices of the form $(v,(m_{1},m_{2}))$ with $v\in V$, $m_{1},m_{2}\in\{0,\ldots,B\}\cup\{\infty\}$, such that, whenever $v$ belongs to Player 00’s target, the weight component $m_{1}$ allows to update the val component $m_{2}$. The initial vertex of $H$ is equal to $(v_{0},0,\textsf{val}(v_{0}))$. This arena has an exponential size and its reachability objective is $\{(v,(m_{1},m_{2}))\mid m_{2}\leq B\}$. It is known that when there is a winning strategy in a zero-sum game with a reachability objective, then there is one that is memoryless [25]. Hence, coming back to $G$, $\sigma_{0}$ can be replaced by a finite-memory strategy with exponential memory.
@@ 5.4 NEXPTIME-Completeness
Let us finally study the NEXPTIME-hardness of the SPS problem. In [16], the Boolean variant of the SPS problem is proved to be NEXPTIME-complete. It can be reduced to its quantitative variant by labeling each edge with a weight equal to 00 and by considering a bound $B$ equal to 00. Hence the val and cost components are either equal to 00 or $\infty$. It follows that the (quantitative) SPS problem is NEXPTIME-hard.
@@ 6 Conclusion and Future Work
In this paper, we suppose that there is one weight function $w:E\rightarrow\mathbb{N}$ used to compute the value and the cost of any play $\rho$, i.e. to compute $\textsf{cost}_{T_{0}}(\rho)$ and $(\textsf{cost}_{T_{1}}(\rho),\ldots,\textsf{cost}_{T_{t}}(\rho))$. We could have considered the more general case where there are $t+1$ weight functions $w_{i}$, $i\in\{0,1,\ldots t\}$, and each cost $\textsf{cost}_{T_{i}}(\rho))$ is computed thanks to $w_{i}$. Unfortunately, the techniques developed in this paper cannot be adapted to this case. Indeed, Lemma 3.3 allows to eliminate cycles, with the restriction that the cycle has a null weight as long as Player 00’s target $T_{0}$ is not visited yet. It follows that in the proof of Proposition 5.1 where we transform each witness into a lasso by eliminating cycles, we can guarantee that this lasso visits $T_{0}$ along its prefix $h$ with length $|V|\cdot B\cdot W$. In the case of games with $t+1$ (dissociated) weight functions $w_{i}$, we do not see how to bound the length of such a prefix $h$.
@@
Considering multiple objectives for Player 00 (instead of one) is also an interesting problem to study. The order on the tuples of values becomes partial. In this case, we could impose a threshold on each value component.

### v3

@@ 4.3 Bounding the Minimum Component of Pareto-Optimal Costs
$\mbox{ with }\quad|C_{\sigma_{0}}|$ $\leq$ $\big{(}f({0},{t})+B+3\big{)}^{t+1}.$ (11)
The proof of Lemma 4.6 requires the next technical property about trees. We recall the notion of depth of a node in a tree: the root has depth 0, and if a node has depth $d$, then its sons have depth $d+1$.
@@
We are going to proceed by induction on $n$. If $n=1$, then the tree only contains nodes of degree at most 1. Thus, the depth of the tree is at most $l-2\leq l\cdot(\log_{2}(n)+1)$.
Let $n\geq 2$ and suppose that the lemma holds for any $k<n$. Consider $\mathcal{T}$ a finite tree with at most $n$ leaves such that there are at most l1l-1 consecutive nodes with degree one along any branch of the tree. Let $X$ be the set of nodes of $\mathcal{T}$ whose degree is greater than or equal to 2. Since $n\geq 2$, $X$ is not empty. Let $x$ be a node in $X$ with the smallest depth. By definition, $x$ is connected to the root of $\mathcal{T}$ by a succession of nodes of degree 1. Thus, its depth is at most l1l-1. Let $d\geq 2$ be the degree of $x$ and $x_{1},\ldots,x_{d}$ be its children. Since $d\geq 2$, there exists $i$ such that the subtree $\mathcal{T}_{i}$ rooted in $x_{i}$ has at most $\lfloor n/2\rfloor$ leaves. Hence, by induction hypothesis, the depth of the leaves of $\mathcal{T}_{i}$ of smallest depth is bounded by $l\cdot(\log_{2}(\lfloor n/2\rfloor)+1)\leq l\cdot\log_{2}(n)$. Since the depth of $x_{i}$ in $\mathcal{T}$ is $l$, the depth of the leaves of $\mathcal{T}$ of smallest depth is bounded by $l+l\cdot\log_{2}(n)=l\cdot(\log_{2}(n)+1)$.
@@
Let us construct the first Pareto-optimal cost $c^{(0)}$. Let $h_{0}$ be the history of maximal length such that $h_{0}\rho$ and $w(h_{0})=B$. This history $h_{0}$ exists because the arena is binary and $\textsf{cost}({\rho})=d$ with $d_{min}$ satisfying (12). Moreover, as $\sigma_{0}$ is a solution, $h_{0}$ visits Player 0’s target. We consider the subtree $\mathcal{T}_{0}$ of $\mathcal{T}$ rooted in the last vertex of $h_{0}$. This subtree has at most $|C_{\sigma_{0}}|$ leaves as it is the case for $\mathcal{T}$. Its nodes with degree $\geq 2$ correspond to histories that are branching points [^9] We recall that a branching point is a history which is the greatest common prefix of two witnesses.. Moreover, any two consecutive nodes with degree one along any branch of ${\mathcal{T}}_{0}$ are in the same region, as Player 0’s target is visited by $h_{0}$ and the first visit to a target of Player $1$ is in a leaf of $\mathcal{T}_{0}$. As $\sigma_{0}$ is without cycle in the sense of Lemma 3.3, it follows that there are at most |V|1|V|-1 consecutive nodes with degree one along any branch of $\mathcal{T}_{0}$.
@@
Let us construct the second Pareto-optimal cost $c^{(1)}$ as we did for $c^{(0)}$. Let $h_{1}$ be the history of maximal length such that $h_{1}\rho$ and $w(h_{1})=B+\delta+1+f({0},{t})$ (notice that we use the bound of (14)). This history $h_{1}$ exists for the same reasons as for $h_{0}$. Let $\mathcal{T}_{1}$ be the subtree rooted in the last vertex of $h_{1}$ (it is a subtree of $\mathcal{T}_{0}$). We can apply Lemma 4.7 as for $\mathcal{T}_{0}$: the leaves of $\mathcal{T}_{1}$ with the smallest depth have a depth $\leq\delta$ . Thus, we set $c^{(1)}$ as the cost of a witness associated with one of these leaves. We get that $c^{(1)}_{min}\leq B+\delta+1+f({0},{t})+\delta$ by definition of $w(h_{1})$. By (7), we get
@@
Finally, we gather all our intermediate results and combine them, in a way to complete the proof of Theorem 4.1. Thanks to Lemmas 4.2-4.6, calculations can be done in a way to have an explicit formula for f(B,t)f({B},{t}) and a bound on its value.
@@
It follows that f(B,t+1)f({B},{t+1}) can be computed thanks to f(0,t)f({0},{t}) bounded by induction:
@@
Therefore, Inequation (4) holds for f(B,t+1)f({B},{t+1}). This completes the proof of Theorem 4.1 by induction on $t$.
@@
In the sequel, we use the same notation f(B,t)f({B},{t}) for any SP games, having in mind that $|V|$ has to be multiplied by $W$ when the game arena is not binary.
@@ 5.1 Finite-Memory Solutions
  • We show that at any deviation [^11] We recall that a deviation is a history $hv$ with $h\in\textsf{Hist}^{1}$, $v\in V$, such that $h$ is prefix of some witness, but $hv$ is prefix of no witness., Player 0 can switch to a punishing strategy that imposes that the consistent plays $\pi$ either satisfy $\textsf{val}(\pi)\leq B$ or $\textsf{cost}({\pi})$ is not Pareto-optimal. Moreover, this punishing strategy is finite-memory with an exponential memory.
@@ 5.2 Punishing Strategies
Let $\sigma_{0}$ be a Pareto-bounded solution to the SPS Problem. By Theorem 4.1 (3) (4), we get that $c_{i}\leq f({B},{t})$ or $c_{i}=\infty$ for all $c\in C_{\sigma_{0}}$ and all $i\in\{1,\ldots,t\}$, such that f(B,t)f({B},{t}) is exponentially bounded. Moreover, if a play $\rho$ is Pareto-optimal, then $\textsf{val}(\rho)\leq B$. We define for each history $g\in\textsf{Hist}_{\sigma_{0}}$ its record $\textsf{rec}({h})=$ $(w(h),\textsf{val}(h),\textsf{cost}({h}))$ whose values are truncated to $\infty$ if they are greater than f(B,t)f({B},{t}).
@@
Let $hv$ be a deviation such that $\textsf{val}(h)=\infty$, that is, $h$ does not visit Player 0’s target. As $\sigma_{0}$ is a solution, notice that $\sigma_{0|hv}$ imposes to each consistent play $\pi$ in $G_{|hv}$ to satisfy $\textsf{val}(\pi)\leq B$ or $\textsf{cost}({\pi})>c$ for some $c\in C_{\sigma_{0}}$. We are going to replace $\sigma_{0|hv}$ by a winning strategy in a zero-sum game $H$ with an exponential arena and an omega-regular objective [^12] We assume that the reader is familiar with the concept of zero-sum game with an omega-regular objective, see e.g. [25]. that is equivalent to what $\sigma_{0|hv}$ imposes to plays. The arena of $H$ is the arena of $G$ extended with the record $\textsf{rec}({g})$ of the current history $g$. More precisely, vertices are of the form $(v,(m_{1},m_{2},m_{3}))$ with $v\in V$, $m_{1},m_{2}\in\{0,\ldots,f({B},{t})\}\cup\{\infty\}$, and $m_{3}\in\big{(}\{0,\ldots,f({B},{t})\}\cup\{\infty\}\big{)}^{t}$, such that, whenever $v$ belongs to some target, the weight component $m_{1}$ allows to update the (truncated) val component $m_{2}$ and the (truncated) cost component $m_{3}$. The initial vertex of $H$ is equal to $(v,\textsf{rec}({hv}))$. This arena is finite and of exponential size by the way the values of $(m_{1},m_{2},m_{3})$ are truncated and by Theorem 4.1 (4). The omega regular objective of $H$ is the disjunction between
@@
Suppose now that $hv$ is a deviation with $\textsf{val}(h)<\infty$. As $h$ already visits Player 0’s target, we can use, in place of $\sigma_{0|hv}$, any memoryless strategy $\tau^{\textsf{Pun}}_{v,\textsf{rec}({hv})}$ as deviating strategy.
@@ 5.3 Lasso Witnesses
(1) We first suppose that $C_{\sigma_{0}}\neq\{(\infty,\ldots,\infty)\}$. According to Theorem 4.1, from $\sigma_{0}\in\mbox{SPS}(G,B)$, one can construct a solution $\sigma_{0}^{\prime}\in\mbox{SPS}(G,B)$ that is Pareto-bounded, thus without cycles. Consider a set of witnesses $\textsf{Wit}_{\sigma^{\prime}_{0}}$ for this strategy, and the region decomposition of its witnesses. Along any witness $\rho$, according to Lemma 3.3, it is thus impossible to eliminate cycles. Recall that Lemma 3.3 only considers cycles implying histories in the same region, those cycles with a null weight before visiting Player 0’s target; moreover, the last region of $\rho$ is excluded. Let us study the memory used by $\sigma^{\prime}_{0}$. Along a witness $\rho$:
  • As $\textsf{val}(\rho)\leq B$ and there is no cycle with a null weight before visiting Player 0’s target, it follows that the smallest prefix $h$ of $\rho$ such that $\textsf{val}(h)=\textsf{val}(\rho)$ has a length $|h|\leq|V|\cdot B\cdot W$ (the presence of $W$ comes from the fact that the arena has been made binary).
  • Once Player 0’s target is visited and before the last region, $\sigma^{\prime}_{0}$ is “locally” memoryless inside each region, as there is no cycle.
@@
  • The number of regions traversed by a witness $\rho$ is bounded by $(t+2)\cdot|\textsf{Wit}_{\sigma^{\prime}_{0}}|$. Indeed, the number of visited targets increases from 0 to $t+1$ and the set $\textsf{Wit}_{\sigma^{\prime}_{0}}(h)$ decreases until being equal to $\{\rho\}$. As there are $|C_{\sigma^{\prime}_{0}}|$ witnesses, the total number of traversed regions is bounded by $(t+2)\cdot|C_{\sigma^{\prime}_{0}}|^{2}$ which is exponential by Corollary 4.11.
@@
(2) Suppose that $C_{\sigma_{0}}=\{(\infty,\ldots,\infty)\}$. This means that each play $\rho\in\textsf{Play}_{\sigma_{0}}$ visits no target of Player $1$ and that $\sigma_{0}$ can impose to each such $\rho$ to have a value $\textsf{val}(\rho)\leq B$. As done in the proof of Lemma 5.2, we can replace $\sigma_{0}$ by a strategy corresponding to a winning strategy in a zero-sum game $H$ with an exponential arena and a reachability objective. The arena of $H$ has vertices of the form $(v,(m_{1},m_{2}))$ with $v\in V$, $m_{1},m_{2}\in\{0,\ldots,B\}\cup\{\infty\}$, such that, whenever $v$ belongs to Player 0’s target, the weight component $m_{1}$ allows to update the val component $m_{2}$. The initial vertex of $H$ is equal to $(v_{0},0,\textsf{val}(v_{0}))$. This arena has an exponential size and its reachability objective is $\{(v,(m_{1},m_{2}))\mid m_{2}\leq B\}$. It is known that when there is a winning strategy in a zero-sum game with a reachability objective, then there is one that is memoryless [25]. Hence, coming back to $G$, $\sigma_{0}$ can be replaced by a finite-memory strategy with exponential memory.
@@ 5.4 NEXPTIME-Completeness
Let us finally study the NEXPTIME-hardness of the SPS problem. In [16], the Boolean variant of the SPS problem is proved to be NEXPTIME-complete. It can be reduced to its quantitative variant by labeling each edge with a weight equal to 0 and by considering a bound $B$ equal to 0. Hence the val and cost components are either equal to 0 or $\infty$. It follows that the (quantitative) SPS problem is NEXPTIME-hard.
@@ 6 Conclusion and Future Work
In this paper, we suppose that there is one weight function $w:E\rightarrow\mathbb{N}$ used to compute the value and the cost of any play $\rho$, i.e. to compute $\textsf{cost}_{T_{0}}(\rho)$ and $(\textsf{cost}_{T_{1}}(\rho),\ldots,\textsf{cost}_{T_{t}}(\rho))$. We could have considered the more general case where there are $t+1$ weight functions $w_{i}$, $i\in\{0,1,\ldots t\}$, and each cost $\textsf{cost}_{T_{i}}(\rho))$ is computed thanks to $w_{i}$. Unfortunately, the techniques developed in this paper cannot be adapted to this case. Indeed, Lemma 3.3 allows to eliminate cycles, with the restriction that the cycle has a null weight as long as Player 0’s target $T_{0}$ is not visited yet. It follows that in the proof of Proposition 5.1 where we transform each witness into a lasso by eliminating cycles, we can guarantee that this lasso visits $T_{0}$ along its prefix $h$ with length $|V|\cdot B\cdot W$. In the case of games with $t+1$ (dissociated) weight functions $w_{i}$, we do not see how to bound the length of such a prefix $h$.
@@
Considering multiple objectives for Player 0 (instead of one) is also an interesting problem to study. The order on the tuples of values becomes partial. In this case, we could impose a threshold on each value component.